### v1

# Hamiltonian Engineering of collective XYZ spin models in an optical cavity: From one-axis twisting to two-axis counter twisting models

Abstract
Quantum simulation using synthetic quantum systems offers unique opportunities to explore open questions in many-body physics and a path for the generation of useful entangled states. Nevertheless, so far many quantum simulators have been fundamentally limited in the models they can mimic. Here, we are able to realize an all-to-all interaction with arbitrary quadratic Hamiltonian or effectively an infinite range tunable Heisenberg XYZ model. This is accomplished by engineering cavity-mediated four-photon interactions between 700 rubidium atoms in which we harness a pair of momentum states as the effective pseudo spin or qubit degree of freedom. Using this capability we realize for the first time the so-called two-axis counter-twisting model, an iconic XYZ collective spin model that can generate spin-squeezed states that saturate the Heisenberg limit bound. The versatility of our platform to include more than two relevant momentum states, combined with the flexibility of the simulated Hamiltonians by adding cavity tones opens rich opportunities for quantum simulation and quantum sensing in matter-wave interferometers and other quantum sensors such as optical clocks and magnetometers.
Introduction
The ability to create and control different many-body interactions is key for entanglement generation, optimization, quantum sensing, and quantum simulation. Long-range interacting systems have recently shown much promise for the engineering of Hamiltonians that generate interesting correlations that propagate across the system. Several experimental platforms are making rapid progress, including Rydberg atoms Browaeys2020 ; Weidemuller2021Floquet ; Wu2021 , polar molecules Bohn2017 ; Ye2023Itinerant , trapped ions Monroe2017TimeCrystal ; Monroe2021 , cavity QED systems Mivehvar2021 and defect centers in solids Lukin2020RobustEngineering ; Munns2023 . Short range contact interactions in ultra-cold atomic systems are another promising approach Bloch2008 ; Schafer2020 ; Gross2017 , however, it remains an open challenge to reach the sufficiently low-temperatures.
So far, most efforts to engineer Hamiltonians have been limited to XXZ spin models or models that feature both exchange and Ising interactions. Common to these models is the fact that the total magnetization of the spin ensemble is preserved. However, limited progress has been achieved in engineering more general spin models, such as XYZ models, which can break both SU(2), and U(1) symmetries and lead to more general ground-state and out-of-equilibrium many-body behaviors. A few exceptions include experiments in Rydberg atoms using two-color dressing Gross2023RydbergXYZ with dynamics limited to pairs of atoms, or Floquet engineering in disordered arrays Weidemuller2021Floquet .
XYZ Hamiltonians can in fact provide important gains in the context of quantum sensing and metrology applications. For instance, the so called two-axis counter-twisting (TACT) KitagawaUedaOAT model is one type of XYZ model well known for its capability to generate entanglement exponentially fast all the way to the Heisenberg bound. It was proposed more than 30 years ago shirasaki1990squeezing ; KitagawaUedaOAT together with a simpler version, the so-called one-axis twisting (OAT) featuring only Ising interactions. So far, only the OAT has been realized in experiments including trapped ions Monroe2000EntanglingFourIons ; Blatt2008IonOAT , Bose-Einstein Condensates OberthalerSqueezedBECInt ; Treutlein2010Squeezing , atomic cavity-QED leroux2010implementation ; VladanEntangledClockLifetime2010 ; norcia2018cavity ; Schleier-Smith2020GapProtection , superconducting qubits Zhu2019SCqubits and optical interferometers AndrianovLeuchs2023KerrSqueezing . Experiments have approximated TACT dynamics locally on the Bloch sphere by combining OAT with a transverse drive (the so-called Lipkin-Meshkov-Glick (LMG) model) Oberthaler2015TaT ; Jessen2007CsTaT ; li2023improving or equivalent approximations in spin-nematic dynamics in higher spin systems ChapmanSqueezing2012 ; You2017TwinFock . However, there has been no demonstration of genuine TACT in any platforms due to the challenge of how to realize this more complex non-linear Hamiltonian.
Here we experimentally show that photon-mediated interactions between atoms inside an optical cavity can be tuned to realize a tunable all-to-all Heisenberg XYZ Hamiltonian, including the TACT model, with the added benefit of not requiring us to prepare extremely low entropy quantum degenerate gases. While there have been prior theory proposals to realize TACT models in atom-cavity systems using four laser tones MonikaSorenson2017NJP ; zhang2017cavity , here we experimentally realize the model using only two dressing lasers. We demonstrate the tunability of the XYZ Hamiltonian by explicitly mapping out the evolution of Bloch vectors on the Bloch sphere at the mean-field level and witness the TACT dynamics for the first time.
The pseudo-spin system here is novel, consisting of two momentum states of atoms freely-falling inside the cavity, making our results of great interest for Bragg matter-wave interferometers LeePRLBragg1995 that are important for both inertial navigation and fundamental science such as searches for dark matter and dark energy, detection of gravitational waves, and determination of the fine structure constant MullerFineStructureAI ; GuellatiKhelifaFineStructure2020 . This approach can also be straightforwardly applied to systems with additional internal levels, making it ideal for developing next-generation quantum-enhanced sensors for technology and exploring a broad range of science from atomic clocks VuleticSqueezedOpticalClock2020 ; robinson2022direct and magnetometers PolzikSqueezedMagnetometer2010 ; MitchellSubSQLMagnetometry2010 ; OberthalerSqueezedMagnetometer2014 to geodesy LudlowGeodesy2018 .
Figure 1: Experimental overview. (A) Illustration of the microscopic momentum pair raising process described by the Hamiltonian $\hat{J}_{+}\hat{J}_{+}$, using two momentum states $p_{0}\pm\hbar k$ as a pseudo-spin 1/2 degree of freedom. Initially, two atoms $i$ and $j$ are in the same momentum state along the cavity axis. Dressing lasers are applied to the cavity (red and blue arrows) that allow atom $i$ to absorb a dressing laser photon and emit a photon (squiggly line) into the cavity such that the net photon recoil flips its momentum state by $2\hbar k$. The emitted photon is absorbed by atom $j$, also flipping its momentum state by $2\hbar k$. There also exists the separate momentum exchange process $\hat{J}_{+}\hat{J}_{-}$, where atoms initially in opposite momentum states flip their momentum states by emitting and absorbing photons explored in Thompson2023MomentumExchange . (B) Frequency diagram of the applied dressing lasers with frequencies $\omega_{1,2}$ and coherent state amplitudes $\alpha_{1,2}$. The emitted photons (squiggly line) are Doppler shifted by $\omega_{z}$ from the dressing laser frequencies and detuned by $\Delta_{c}$ from the cavity resonance frequency $\omega_{c}$. The cavity is far detuned from the atomic transition frequency $\omega_{a}$. (C) Representation of the emission/absorption processes described in (A) but depicted in terms of the atomic energy versus momentum for atoms $i$ and $j$. The reverse process is also allowed, giving rise to a collective lowering operator described by $\hat{J}_{-}\hat{J}_{-}$. (D) Dynamics induced by two-axis counter twisting in the form of a Hamiltonian $\hat{J}_{x}^{2}-\hat{J}_{z}^{2}$ is represented on a Bloch sphere with north and south poles defined by |$\ket{p_{0}\pm\hbar k}$. Top right shows the local circular flows around the stable point at $-\hat{x}$. The local flows for unstable saddle points at $\hat{y}$ with exponential squeezing and anti-squeezing along $\hat{x}\pm\hat{z}$ are shown in bottom right. (E) Observation of the four-photon resonance that generates pair raising and lowering processes. We scan the dressing laser frequency difference to vary the detuning $\delta$. With equal dressing laser amplitudes corresponding to realizing the $\hat{J}_{x}^{2}$ Hamiltonian when $\delta=0$, we see clear resonances in the observed change of the Bloch vector’s azimuthal angle $\phi$ for a Bloch vector prepared near the south pole ($\theta=\pi/4$, red data points) and near the north pole ($\theta=3\pi/4$, blue data points). All error bars reported are $1\sigma$ uncertainties. Simulation results are shown in solid lines with shaded area allowing for 5% uncertainty in interaction strength. See Methods for details.
Experimental setup
In the experiment, ${}^{87}$b atoms are laser-cooled inside a vertically-oriented two-mirror standing wave cavity, see Fig. 1(A) and Thompson2023MomentumExchange ; Thompson2022SAI . A repulsive intra-cavity doughnut dipole trap confines the atoms radially, but allows the atoms to fall along the cavity axis. To prepare atoms in a well-defined momentum state, a pair of laser beams are injected into the cavity to drive velocity-dependent two-photon Raman transitions between ground hyperfine states |$\ket{F=1,m_{F}=0}$ and |$\ket{F=2,m_{F}=0}$. After removing the unselected atoms with a resonant laser push beam, successive microwave pulses are applied to prepare the internal states of the selected atoms in the ground hyperfine state |$\ket{F=2,m_{F}=2}$ (see Methods and Thompson2023MomentumExchange ).
With about 700 atoms centered at momentum $p_{0}-\hbar k$, another pair of laser beams is injected along the cavity axis to drive two-photon Bragg transitions connecting the two momentum states |$\ket{p_{0}\pm\hbar k}$ which defines a two-level spin-1/2 system Thompson2022SAI ; Thompson2023MomentumExchange . Here the average momentum is $p_{0}$, $\hbar$ is the reduced Plank constant, and $k=2\pi/\lambda$ where $\lambda$ is the wavelength of the Bragg laser beams. Due to finite momentum spread, one can consider momentum wave packets centered at these two momentum states as considered in detail previously Thompson2023MomentumExchange . Ignoring the finite momentum spread of the selected momentum states, we define $\hat{\psi}_{\uparrow,\downarrow}^{\dagger}$ and $\hat{\psi}_{\uparrow,\downarrow}$ as the operators for creating and annihilating an atom in momentum states ||$\ket{\uparrow}\equiv\ket{p_{0}+\hbar k}$ and ||$\ket{\downarrow}\equiv\ket{p_{0}-\hbar k}$. For mapping to a pseudo-spin model, we define ladder operators $\hat{J}_{+}=\hat{\psi}_{\uparrow}^{\dagger}\hat{\psi}_{\downarrow}$, $\hat{J}_{-}=\hat{\psi}_{\downarrow}^{\dagger}\hat{\psi}_{\uparrow}$ and spin projection operators $\hat{J}_{x}=\frac{1}{2}\left(\hat{J}_{+}+\hat{J}_{-}\right)$, $\hat{J}_{y}=\frac{1}{2i}\left(\hat{J}_{+}-\hat{J}_{-}\right)$ and $\hat{J}_{z}=\frac{1}{2}\left(\hat{\psi}_{\uparrow}^{\dagger}\hat{\psi}_{\uparrow}-\hat{\psi}_{\downarrow}^{\dagger}\hat{\psi}_{\downarrow}\right)$.
As shown in Fig. 1(B), the cavity frequency $\omega_{c}$ is detuned from the atomic transition |$\ket{F=2,m_{F}=2}\rightarrow\ket{F^{\prime}=3,m_{F^{\prime}}=3}$ by $\Delta_{a}=\omega_{c}-\omega_{a}=2\pi\times 500$ MHz, which is much larger than the excited state decay rate $\Gamma=2\pi\times 6~{}$MHz and the cavity power decay rate $\kappa=2\pi\times 56(3)~{}$kHz. A series of Bragg pulses can be applied to realize a Mach-Zehnder matter-wave interferometer (i.e. $\pi/2-\pi-\pi/2$), in which the wave packets first separate in position and then re-overlap. When the two wave packets are overlapped, the interference between them forms an atomic density grating with period $\lambda/2$, which matches the standing wave of a cavity mode.
As the atoms move along the cavity axis, the density grating is periodically aligned to the standing-wave of the cavity mode, leading to modulation of the cavity resonance frequency at the two-photon Doppler frequency $\omega_{z}=2kp_{0}/m\approx 2\pi\times 500~{}$kHz, where $m$ is the mass of ${}^{87}$b. To have this modulation to mediate an effective atom-atom interaction, we typically apply two $\sigma^{+}$ polarized dressing laser tones (see Fig. 1(A)) at frequencies $\omega_{1,2}$ within a few MHz of the cavity resonance frequency (see Fig. 1(B)), with complex amplitudes $\alpha_{1,2}$ corresponding to the field that would be established inside the cavity were no atoms in the cavity and implicit units $\sqrt{\mathrm{photons}}$. The atom–induced cavity frequency modulation leads to the generation of modulation sideband tones at frequencies $\omega_{1,2}\pm\omega_{z}.$ In the following simplifications, we will assume $\omega_{2}>\omega_{1}$.
The key insight is that different combinations of the dressing lasers and their atom-induced sideband tones will induce different virtual four-photon processes which will manifest as all-to-all exchange interactions $\hat{J}_{+}\hat{J}_{-}$ Thompson2023MomentumExchange and pair-raising $\hat{J}_{+}\hat{J}_{+}$ (lowering $\hat{J}_{-}\hat{J}_{-}$) processes as shown in Fig. 1(A,C). After adiabatically eliminating the cavity fields using second order perturbation theory (see Thompson2023MomentumExchange and Methods), we obtain an effective time-dependent atom-only Hamiltonian in an appropriate frame rotating at $\omega_{z}$
$\hat{H}=\chi_{e}\hat{J}_{+}\hat{J}_{-}+\left(\chi_{p}e^{i\delta t}\hat{J}_{+}\hat{J}_{+}+\chi_{p}^{*}e^{-i\delta t}\hat{J}_{-}\hat{J}_{-}\right),$ (1)
with the exchange and pair-raising/lowering couplings given by
$\begin{split}\chi_{e}&=\left(\frac{g_{0}^{2}}{4\Delta_{a}}\right)^{2}\left(\frac{|\alpha_{1}|^{2}}{\Delta_{\bar{c}}+\delta/2}+\frac{|\alpha_{2}|^{2}}{\Delta_{\bar{c}}-\delta/2}\right),\\ \chi_{p}&=\left(\frac{g_{0}^{2}}{4\Delta_{a}}\right)^{2}\frac{|\alpha_{1}\alpha_{2}|e^{i\phi_{\mathrm{int}}}}{2}\left(\frac{1}{\Delta_{\bar{c}}+\delta/2}+\frac{1}{\Delta_{\bar{c}}-\delta/2}\right).\end{split}$ (2)
Here, $\Delta_{\bar{c}}=\left(\omega_{2}+\omega_{1}\right)/2-\omega_{c}$ is the average detuning of the two dressing lasers from cavity resonance, typically set to be less than 1 MHz. $\delta=\left(\omega_{2}-\omega_{1}\right)-2\omega_{z}$ is the detuning from four-photon resonance, $g_{0}=2\pi\times 0.48$ MHz is the maximal atom-cavity Rabi coupling at an anti-node of the cavity mode, and $\phi_{\mathrm{int}}=\arg\left(\alpha_{2}\alpha_{1}^{*}\right)-\phi_{\mathrm{B}}$ is the differential phase between the two dressing laser tones relative to the phase of Bragg coupling $\phi_{\mathrm{B}}$ which forms initial density grating (see Methods). Collective cavity dissipation is dealt with separately (see Methods).
We will focus on the resonant case $\delta=0$ and $\phi_{\mathrm{int}}=0$, though we show example data in Fig. 1(E) (see Method) that clearly exhibits a resonance in the interaction-induced dynamics at $\delta=0$ as the pair raising/lowering processes are tuned into and out of resonance by tuning the dressing laser frequency difference. In the resonant case, the Hamiltonian of Eq. (1) reduces to:
$\hat{H}$ $=\left(\chi_{e}+\chi_{p}\right)\hat{J}^{2}_{x}+\left(\chi_{e}-\chi_{p}\right)\hat{J}^{2}_{y}$ (3)
$=\chi_{e}\hat{\bf J}\cdot\hat{\bf J}+\chi_{p}\left(\hat{J}_{x}^{2}-\hat{J}_{y}^{2}\right)-\chi_{e}\hat{J}_{z}^{2},$
where we defined the collective angular momentum operator, $\hat{\bf J}=\{\hat{J}_{x},\hat{J}_{y},\hat{J}_{z}\}$ and introduced the collective Heisenberg interaction $\hat{\bf J}\cdot\hat{\bf J}$. The latter acts as a constant for any eigenstate of the collective angular momentum operator $\hat{\bf J}\cdot\hat{\bf J}$, such as the collective states with eigenvalue $N/2(N/2+1)$. In the presence of single particle inhomogeneities, this term opens a many-body gap that help promote spin locking which we explored before for the protection of coherences against dephasing norcia2018cavity ; Thompson2023MomentumExchange . Here we instead focus on the fully collective dynamics and therefore without loss of generality we can add a generic $\chi_{z}\hat{\bf J}\cdot\hat{\bf J}$ term without affecting the dynamics. As such, in our system we are able to engineer dynamics governed by an XYZ Hamiltonian $\hat{H}=\chi_{x}\hat{J}^{2}_{x}+\chi_{y}\hat{J}^{2}_{y}+\chi_{z}\hat{J}^{2}_{z}$ with interaction strengths $\chi_{x}=\left(\chi_{e}+\chi_{p}+\chi_{z}\right)$ and $\chi_{y}=\left(\chi_{e}-\chi_{p}+\chi_{z}\right)$. The XYZ Hamiltonian is highly tunable by simply adjusting the relative power in the two applied dressing lasers since $\chi_{e}$ scales as $\left|\alpha_{2}\right|^{2}$, $\left|\alpha_{1}\right|^{2}$ and $\chi_{p}$ scales as $\left|\alpha_{2}\alpha_{1}\right|$.
Figure 2: Evolution under different Hamiltonians (A) Setting different amplitude ratios between the two dressing tones $|\alpha_{2}/\alpha_{1}|$ gives rise to different XYZ Hamiltonians, with specific examples shown here including one-axis twisting $\hat{\bf J}\cdot\hat{\bf J}-\hat{J}_{z}^{2}$, two-axis counter twisting $\hat{\bf J}\cdot\hat{\bf J}+\hat{J}_{x}^{2}-\hat{J}_{z}^{2}$ and one-axis twisting $\hat{J}_{x}^{2}$. The experimentally observed dynamics are shown in corresponding rows in (B, C, D), where the tail of each vector indicates the initial position of the Bloch vector $\mathbf{J}_{i}$ on the Bloch sphere, and the arrow indicates the displacement $\mathbf{T}$ after a brief period of evolution under the corresponding Hamiltonians. The left (right) panels are for initial Bloch vectors on the south (north) hemisphere. The polar angle $\theta$ of the initial Bloch vector linearly increase from $\pi/2$ at the rim to $\pi$ in the middle for the left plot, whereas $\theta$ decrease from $\pi/2$ at the rim to 00 in the middle for the right plot. The middle panels are equirectangular projections. In each case, the qualitatively observed stable fixed points are marked with numbered red circles, and the unstable fixed points are marked with numbered red squares. The blue lines are shown to indicate directions of observed flow. The red lines indicate lines where the dynamics are zero and separate regions of opposite flows on the Bloch sphere.
Mean-field dynamics
At the mean-field level, we can define the Bloch vector $\mathbf{J}\equiv(J_{x},J_{y},J_{z})=(\langle\hat{J}_{x}\rangle,\langle\hat{J}_{y}\rangle,\langle\hat{J}_{z}\rangle)$ and approximate the Hamiltonian as $\hat{H}=\mathbf{B}(\mathbf{J})\cdot\hat{\mathbf{J}}$. In this way, the collective dynamics are driven by a self-generated effective magnetic field $\mathbf{B}(\mathbf{J})=(2\chi_{x}{J}_{x},2\chi_{y}{J}_{y},2\chi_{z}{J}_{z})$ which depends on the instantaneous collective spin projections.
We can derive the equations of motion of the collective Bloch vector, from Eq. (3) (see Method), which simplify to a nonlinear torque equation $d\mathbf{J}/dt=\mathbf{B}(\mathbf{J})\times\mathbf{J}\equiv\mathbf{T}(\mathbf{J})$. One can identify the fixed points $\mathbf{J}_{\rm fix}$ as the points where $\mathbf{T}(\mathbf{J}_{\rm fix})=0$.
To understand the dynamics near the fixed points, it is useful to follow a standard stability analysis by diagonalizing the Jacobian matrix $M(\mathbf{J})=\partial\mathbf{T}/\partial\mathbf{J}|_{\mathbf{J}=\mathbf{J}_{\rm fix}}$. The local motion near these fixed points is illustrated in Fig. 1(D) (right). We use red circles for stable points with purely imaginary eigenvalues. The Bloch vector evolves on stable closed orbits indicated by the blue circular traces. The red squares are used to denote unstable saddle points with real eigenvalues with opposite signs. The unstable saddle points are labelled by the red squares. The eigenvalues of the Jacobian matrix at the saddle points are real but with opposite signs. In Fig. 1(D) (bottom right), the dynamics show exponential divergence from the origin (indicated by outward blue arrows) along $\hat{x}+\hat{z}$, corresponding to the positive eigenvalue PoggiDeutsch2023PRXQuantum . The negative eigenvalues are indicated by the convergence towards the origin (inward blue arrows) along $\hat{x}-\hat{z}$.
Dynamics on the Bloch sphere
In the experiment, we probe the local dynamics induced by the above Hamiltonian with various values of $\chi_{e}$ and $\chi_{p}$. To do this, we vary the phase and duration of the Bragg pulse to prepare an initial pseudo-spin coherent state $\mathbf{J}_{i}$. Before the atomic wave packets separate, we then apply the interaction for a short time $\Delta t$ satisfying $\chi N\Delta t\ll 1$ and measure the change in azimuthal angle $d\phi$ and polar angle $d\theta$ to obtain the final Bloch vector $\mathbf{J}_{f}$ after the interaction. This is achieved by repeating the experiment and applying additional appropriate rotations before measuring populations in the two momentum states (see Methods). The local flow vector is then determined by the torque $\mathbf{T}(\mathbf{J}_{i})\approx\Delta\mathbf{J}/\Delta t=\left(\mathbf{J}_{f}-\mathbf{J}_{i}\right)/\Delta t$.
In Fig. 2(A), we show the predicted flow vectors $\mathbf{T}(\mathbf{J}_{i})$ on the Bloch sphere for three example Hamiltonians of interest. Different Hamiltonians are obtained by changing the ratio of the dressing laser amplitudes $\left|\alpha_{2}/\alpha_{1}\right|=0,0.17$, and $1.0$.
We show the measured flow vector $\mathbf{T}(\mathbf{J}_{i})$ in Fig. 2(B) with each row aligned to the corresponding example presented in Fig. 2(A). In each case, the flow vector start at $\mathbf{J}_{i}$ and end at $\mathbf{J}_{f}$. The left and right panels are polar plots (radial coordinate linear in polar angle) of the dynamics on the south/north hemisphere looking from the north poles of the Bloch sphere. The middle panels are equirectangular projections to show dynamics near the equator. From these vector maps, we can make qualitative comparisons based on the geometry of the flow. For both theoretical and experimental results, the stable fixed points and unstable saddle points labeled as numbered red circles and squares respectively PoggiDeutsch2023PRXQuantum ; Witkowska2015TACTdynamics .
In the first row of Fig. 2, we consider the simplest case with $\chi_{p}=0$. This is achieved by turning off one of the dressing lasers ($\left|\alpha_{2}/\alpha_{1}\right|=0$), leading to the Hamiltonian $\hat{H}=\chi_{e}\left(\hat{\bf J}\cdot\hat{\bf J}-\hat{J}_{z}^{2}\right)$. This Hamiltonian, referred to as the OAT Hamiltonian KitagawaUedaOAT , maintains U(1) symmetry, thereby conserving ${J}_{z}$. At the mean-field level, it features a constant effective magnetic field along the $\hat{z}$ direction, which results in the rotation of the collective Bloch vector about the $\hat{z}$-axis at a uniform constant angular frequency,$-2\chi_{e}{J}_{z}$. As expected, we observe two stable fixed points (red 1 and 2 circles), and a reversal of the circulation across the equator (i.e. $J_{z}=0$) where there are no dynamics (red line). This $J_{z}$ dependent circulation leads to shearing of the quantum noise in the orientation of a Bloch vector on the Bloch sphere, a semiclassical explanation for how OAT dynamics generate spin-squeezed states KitagawaUedaOAT ; Thompson2022SAI . We note that while the term $\hat{\bf J}\cdot\hat{\bf J}$ is trivial for our current observations at short times, at longer times, when inhomogeneities in our system manifest, it can lead to important dynamical effects as shown in Ref. Thompson2023MomentumExchange .
Next, we consider the last row in Fig. 2 with $\chi_{p}=\chi_{e}$. This is achieved by using equal dressing laser amplitudes $\left|\alpha_{2}/\alpha_{1}\right|=1$. Here, the Hamiltonian is $\hat{H}=2\chi_{e}\hat{J}_{x}^{2}$, leading to OAT dynamics along the $\hat{x}$ direction. The corresponding dynamics are induced at the mean-field level by a magnetic field along $\hat{x}$ that preserves ${J}_{x}$ and induces a rotation about $\hat{x}$ with constant angular frequency $4\chi_{e}{J}_{x}$. It is noteworthy that the interaction strength here is twice that of the $\hat{J}^{2}_{z}$ case, attributed to the use of two dressing laser tones. The data qualitatively shows two stable fixed points along $\hat{x}$ labeled by red circles 1 and 2, and a reversal of the sign of circulation across ${J}_{x}$ highlighted by the red lines.
Finally, we come to the case that achieves TACT, as shown in the middle row of Fig. 2. In this case, the ratio of dressing lasers amplitudes is set to be $|\alpha_{2}/\alpha_{1}|=\left(\sqrt{2}-1\right)/\left(\sqrt{2}+1\right)\approx 0.17$, which produces $\chi_{e}=3\chi_{p}$. The Hamiltonian then become $\hat{H}=2\chi_{p}\left(2\hat{J}_{x}^{2}+\hat{J}_{y}^{2}\right)=2\chi_{p}\left(\hat{\bf J}\cdot\hat{\bf J}+\hat{J}_{x}^{2}-\hat{J}_{z}^{2}\right)$.
The corresponding mean-field magnetic field is now $\mathbf{B}(\mathbf{J})=4\chi_{p}\left({J}_{x},0,-{J}_{z}\right)$, where we have again ignored the Heisenberg term since we will only consider dynamics at constant Bloch vector length. The theoretical flow line for the dynamics is depicted in Fig. 2(A) (middle), showing four stable fixed points along ±$\pm\hat{x}$ and ±$\pm\hat{z}$, as well as two unstable fixed points along ±$\pm\hat{y}$ connected by great circles inclined at ±$\pm\pi/4$ to the equatorial plane (red circles). In comparison to OAT, there exist two twisting fields thus only the two points intersected by two blue circles with ${J}_{x}={J}_{z}=0$, exhibit the maximum shearing dynamics, which serve as the unstable saddle points.
We experimentally explore the flow lines. In the left and right panels, the observed stable fixed points are labeled with red circles numbered 1 to 4. The circulations of the flow lines are opposite for stable fixed points on opposite sides of the Bloch sphere. In the middle panel, we observe two unstable points labeled with red squared 1 and 2 in Fig. 2(C). The two unstable fixed points have flow lines that either diverge from them (blue arrows outward) or converge towards them (blue arrows inward), with the two flows orthogonal to each other (see the discussion below). These findings constitute the first direct observation of genuine TACT dynamics. We note that for a particular set of parameters of the LMG Hamiltonian, for example when $\hat{H}=\chi\hat{J}_{z}^{2}+\delta\hat{J}_{y}$ with $\delta\sim\chi N$, and when the Bloch vector initially points along the $\hat{y}$ direction PoggiDeutsch2023PRXQuantum ; Witkowska2015TACTdynamics , the flow lines can resemble the ones close to a saddle point of the TACT model (see Methods). However, the instability is restricted to this single point, in contrast to the full TACT which features two independent unstable points.
Figure 3: Dynamics near saddle points. (A) Measured local flow vector map around the saddle point of the TACT dynamics. (B) Blue and red points are parallel projections of the local flow vector onto the $\hat{n}_{\pm}$ axis $\Delta\mathbf{J}\cdot\hat{n}_{\pm}/\left(N/2\right)$ as a function of initial displacement $d\mathbf{J}_{i}\cdot\hat{n}_{\pm}/\left(N/2\right)$ from the saddle point along the blue and red lines in (A). Blue and red solid lines are simulation results. The linear dependence shows exponentially growing / shrinking dynamics along the squeezing / anti-squeezing axes. (C) Green and orange points are the perpendicular projection of the local flow vector as a function of initial angular displacement from the saddle point along the green and orange lines in (A) ($\hat{n}=\hat{x}$ and $\hat{z}$) onto $\hat{n}\times\hat{y}=\hat{z}$ and $-\hat{x}$ axis ($\Delta\mathbf{J}_{z}$ and $-\Delta\mathbf{J}_{x}$). Orange and green solid lines are simulation results.
For a quantitative comparison, we examine the dynamics near the saddle point $\mathbf{J}_{\rm sad}$ along the $\hat{y}$ axis on the Bloch sphere. In Fig. 3(A), we map the displacement $\mathbf{T}\left(\mathbf{J}_{i}\right)$ as a function of the initial Bloch vector orientation $\mathbf{J}_{i}=\mathbf{J}_{\rm sad}+d\mathbf{J}_{i}$. We scan the initial Bloch vector angles $\theta_{i}$ and $\phi_{i}$ over a range ±$\pm\pi/12$ centered about $\pi/2$ (i.e. about $\hat{y}$) with discrete points sampled using a detailed $11\times 1$ 11 grid. The mean-field equations of motion for the two orthogonal directions $\hat{n}_{\pm}=\left(\hat{x}\pm\hat{z}\right)/\sqrt{2}$ are:
$\frac{d}{dt}({J}_{x}+{J}_{z})$ $=4\chi_{p}{J}_{y}({J}_{x}+{J}_{z})$ (4)
$\frac{d}{dt}({J}_{x}-{J}_{z})$ $=-4\chi_{p}{J}_{y}({J}_{x}-{J}_{z}).$
For the small range of angles around the $\hat{y}$ axis sampled in these measurements, one can assume ${J}_{y}\approx N/2$, and therefore find out that the displacement’s time derivative increases linearly with the displacement itself, indicating dynamics that change exponentially over time.
We first focus on the data points along the two directions $\hat{n}_{\pm}$ (depicted by blue and red lines in Fig. 3(A)) and compute the parallel projection of the flow $\Delta\mathbf{J}\cdot\hat{n}_{\pm}/\left(N/2\right)$ as a function of initial displacement from the saddle point $d\mathbf{J}_{i}\cdot\hat{n}_{\pm}/\left(N/2\right)$, as shown in Fig. 3(B). The observed linear relationship between the projection of the flow and initial displacement matches well with the prediction from the simulation results (solid lines) which go beyond the linear approximation by solving the non-linear equations (Eq. (4)). The small differences observed between the $\Delta\mathbf{J}\cdot\hat{n}_{\pm}/\left(N/2\right)$ slopes stem from the finite duration of the interaction, which extends the dynamics beyond the linear response regime.
The unstable dynamics explain why an initial circular distribution centered around the saddle point shears as a function of time by squeezing (anti-squeezing) along the $\hat{n}_{-}$ ($\hat{n}_{+}$) direction at a rate exponentially faster than the linear growth seen in OAT. This combined with the global (over the full Bloch sphere) dynamical behavior allows the TACT to directly approach the fundamental Heisenberg limit on phase estimation KitagawaUedaOAT ; PoggiDeutsch2023PRXQuantum .
In Fig. 3(C), we analyze the orthogonal projection of interaction-induced flows, especially focus on the data points that initially displaced from the saddle point along $\hat{n}=\hat{x}$ and $\hat{z}$ axes (depicted by green and orange lines in Fig. 3(A)), and calculate their projections along the $\hat{n}\times\hat{y}=\hat{z}$ and $-\hat{x}$ axes ($\Delta\mathbf{J}_{z}$ and $-\Delta\mathbf{J}_{x}$), respectively. The dynamics can be explained by noticing that when the Bloch vector is initially prepared in the $y$-$z$ plane, the effective mean-field magnetic field is along the $\hat{z}$ axis with a magnitude of $-4\chi_{p}{J}_{z}$. Conversely, when prepared in the $y$-$x$ plane, the field is along the $\hat{x}$ axis, with an amplitude of $4\chi_{p}{J}_{x}$. Thus one expects these two perpendicular displacements to grow linearly in magnitude with initial Bloch vector displacement, as we observed in Fig. 3(C).
Two-axis counter-twisting with unstable points at north and south poles
The original TACT Hamiltonian, as proposed by Kitagawa and Ueda KitagawaUedaOAT , is defined as
$\hat{H}^{\prime}=\chi\left(\hat{J}_{+}^{2}+\hat{J}_{-}^{2}\right)=2\chi\left(\hat{J}_{x}^{2}-\hat{J}_{y}^{2}\right),$ (5)
with the theoretical flow lines shown in Fig. 4(A). This Hamiltonian bears resemblance to the previously discussed TACT, with an additional $\pi/2$ rotation around the $y$-axis. It is characterized by unstable points at the north and south poles of the Bloch sphere, along with four stable points located on the equator at ±$\pm\hat{x}$ and ±$\pm\hat{y}$. In our setup, it seems to be impossible to choose the parameters canceling $\chi_{e}$ in Eq. (2) at first glance. However, this can be achieved by placing the two detuned dressing lasers tones on the same side of the cavity resonance as shown in Fig. 4(B) (top). In this configuration, the previously ignored lower modulation sideband of the red dressing laser becomes significant. This sideband introduces an exchange interaction with the opposite sign of the other generated exchange interactions. By carefully selecting the detuning and the amplitude ratio of the two dressing laser tones, we can achieve a configuration where $\chi_{e}=0$, effectively only leaving the $\chi_{p}$ term (see Methods.) $\hat{H}^{\prime}$ offers the advantage of possibly being less sensitive to the presence of superradiance or collective decay for future squeezing generation in our system.
Figure 4: Two-axis counter-twisting with unstable points at north/south poles. (A) Dynamics shown as flows on the Bloch sphere with three characteristic circles (purple: $\theta_{i}=0.9\pi$, black: $\theta_{i}=0.5\pi$ and green: $\theta_{i}=0.1\pi$) highlighted. (B) Frequency diagram with both dressing lasers positioned on the same side relative to the cavity resonance (top). The equirectangular projection of the resulting dynamics (bottom). (C), (D) and (E) The distribution of the states after the interaction with the prepared initial state $\theta_{i}=0.1\pi$ (green circle), $0.9\pi$ (purple circle) and $0.5\pi$ (black circle). The deviations from the initial state distribution (gray dashed circles) are shown by the colored arrows.
In Fig. 4(B) (bottom), we show the measured flow lines for $\hat{H}^{\prime}$ in the equirectangular projection. One can clearly identify four stable fixed points on the equator as expected. To have a better intuition of the dynamics, instead of focusing on the dynamics on the whole Bloch sphere, we take a few cuts with initial Bloch vectors $\mathbf{J}_{i}$ prepared with $\theta_{i}=0.1\pi,0.5\pi$ and $0.9\pi$ and study the dynamics separately. For the initial Bloch vectors prepared on the two circles with $\theta_{i}=0.1\pi$ and $0.9\pi$ (green and purple in Fig. 4(A) and (C)/(D) near the north/south poles, the distribution of the states after the interaction $\mathbf{J}_{f}$ is plotted in Fig. 4(C) and (D) (solid lines with fitted curves). The elliptical distributions with major axis orthogonal to each other explicitly show the squeezing and anti-squeezing axis near the north and south poles in this small displacement limit. In Fig. 5(E), the initial Bloch vector is prepared on the equator with different azimuthal angles (grey dashed line) and we plot $\mathbf{J}_{f}$ (black dot). The four zero crossings correspond to the four stable points on the equator. Between the stable points, the observed final states are deflected alternately above or below the equator as expected.
Conclusion
Here we have demonstrated the flexibility of our optical cavity simulator to engineer tunable XYZ Hamiltonians using two selected momentum states, without the need for Floquet engineering which in some cases might be challenging to allow access beyond mean-field dynamics in large many-body systems due to amplitude and phase noise on the Floquet control fields ThompsonPhaseNoise2012 . The pair-raising/lowering processes are verified to be present for the first time by the resonant spectroscopic signal when the detuning $\delta$ is scanned as well as the direct cancellation of the exchange interactions that yielded dynamics of the Hamiltonian $\hat{H}^{\prime}=\chi\left(\hat{J}_{+}^{2}+\hat{J}_{-}^{2}\right)$. By combining the correct relative balance of exchange and pair-raising/lowering contributions, we observed TACT dynamics for the first time.
While so far our focus has been on the characterization of the Hamiltonian by probing the short-time dynamics, extension to longer times can open untapped opportunities for quantum simulation and metrology. For example, by using an implementation where two dressing lasers are used to generate one-axis twisting dynamics, one should be able to balance the collective dissipation generated by the absorbed and emitted photons facilitating the emergence of beyond mean-field effects and entanglement generation.
Furthermore, while here we focused on only two momentum states as a first step, by combining them with actual internal states of the atoms or by adding more selected momentum states and dressing tones one should be able to engineer a toolbox for quantum state engineering, similar to what has been done with momentum states in Bose-Einstein condensate meier2016observation ; An2018 ; Meir2018 ; Gadway2021MomentumLattice . In our case however, in addition to the internal and external level control that tunes the synthetic dimensions, we can further enjoy the rich opportunities offered by the tunable cavity-mediated interactions to engineer phenomena ranging from superfluidity and supersolidity baumann2010dicke ; Julian2017 ; leonard2017monitoring ; schuster2020supersolid to dynamical gauge fields and non-trivial topological behaviors DynamicalPotential2013RMP ; EsslingerSpinTexture2018 ; Esslinger2022DynamicalCurrent ; Mivehvar2021 ; Ritsch2022Dynamical .